# Open Heavy flavour mesons in hot asymmetric strange hadronic matter – a QCD sum rule approach

### v1

@@ IV Results and discussions
In fig. 11, the Borel curves for the strange-bottom mesons ($B_{s}^{0}$ and $\bar{B_{s}}^{0})$ are shown for density $\rho_{B}=2\rho_{0}$ and Fig. 12 shows the density dependence of these mesons for the nuclear and the strange hadronic matter. The mass of the antiparticle (${B_{s}^{0}}$) is observed to have lower mass as compared to the mass of the particle ($\bar{B_{s}^{0}}$) in nuclear matter ($f_{s}$=0) for which $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$ is zero, whereas, the opposite trend is observed when there is non-zero strangeness in the medium. The lower (higher) mass of the particle ($\bar{B_{s}^{0}}$) as compared to the antiparticle ($B_{s}^{0}$) in the nuclear (hyperonic) matter shows that the term $\langle s^{\dagger}s\rangle$ dominates over the other two terms in the odd part of the spectral function, which is responsible for the splitting of the masses between the particle and antiparticle. The behaviour of the mass splitting at finite densities is similar to the case of strange-charm mesons (as shown in Figs. 5 and 6). Similar to the strange-charm mesons ($D_{s}^{\pm}$), the isospin asymmetry leads to a drop (rise) in the mass of $\bar{B_{s}^{0}}({B_{s}^{0}})$ for nuclear (hyperonic) matter. However, the mass modifications due to isospin asymmetry are observed to be marginal for T=50 and 100 MeV in hyperonic matter.
@@ V Summary
In the present work, we have studied the medium modifications of the masses of the open charm ($D$, $\bar{D}$, $D_{s}$ and $\bar{D}_{s}$) and open bottom ($B$, $\bar{B}$, $B_{s}$ and $\bar{B}_{s}$) in isospin asymmetric strange hadronic matter at finite temperature. These are computed using a QCD sum rule approach retaining the QCD operators upto dimension 5 in the operator product exapnsion. The values of the QCD operators are calculated within a chiral SU(3) model. The mass splitting of the particles ($D^{0}$, $D^{+}$, $D_{s}^{+}$, $B^{-}$, $\bar{B^{0}}$, $\bar{B_{s}^{0}}$) and their antiparticles ($\bar{D^{0}}$, $D^{-}$, $D_{s}^{-}$, $B^{+}$, ${B^{0}}$, ${B_{s}^{0}}$) arises due to the odd part of the spectral function. In nuclear matter, the masses of the particles, $D^{+}(D^{0})$ and $B^{-}(\bar{B^{0}})$ mesons are observed to be smaller than the masses of the antiparticles, $D^{-}({\bar{D}}^{0})$ and $B^{+}({B^{0}})$. For the strange-charm as well as strange-bottom mesons, the opposite behaviour of the masses of particles ($D_{s}^{+}$ and $\bar{B_{s}^{0}}$) to be higher than the masses of antiparticles ($D_{s}^{-}$ and $B_{s}^{0}$) is observed in nuclear matter, whereas, in the strange hadronic matter, the non-zero value of $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$ leads to the masses of the antiparticles ($D_{s}^{-}$ and $B_{s}^{0}$) to be higher than that of $D_{s}^{+}$ and $\bar{B_{s}^{0}}$ in hyperonic medium. The sign of the difference in the masses of the particle and antiparticle ($m_{+}-m_{-}$) is observed to be opposite in nuclear and hyperonic matter, both for the strange-charm as well as strange-bottom mesons, In the mass splitting between the particle and antiparticle arising due to the odd part of the spectral function, the term $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$ for the strange-charm (strange-bottom) mesons thus dominates over the other QCD operators $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$. The effects from density, isospin asymmetry, strangeness as well as temperature on the masses of the open charm (bottom) mesons are observed to be appreciable. These can modify e.g., the production of the charm mesons in asymmetric relativistic heavy ion collision experiments planned at the future facility at GSI, as well as in experiments at J-PARC.

### v2

@@ IV Results and discussions
In fig. 11, the Borel curves for the strange-bottom mesons ($B_{s}^{0}$ and $\bar{B_{s}}^{0})$ are shown for density $\rho_{B}=2\rho_{0}$ and Fig. 12 shows the density dependence of the mass shifts of these mesons for the nuclear and the strange hadronic matter. The mass of the antiparticle (${B_{s}^{0}}$) is observed to have lower mass as compared to the mass of the particle ($\bar{B_{s}^{0}}$) in nuclear matter ($f_{s}$=0) for which $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$ is zero, whereas, the opposite trend is observed when there is non-zero strangeness in the medium. The lower (higher) mass of the particle, $\bar{B_{s}^{0}}$ as compared to the antiparticle, $B_{s}^{0}$ in the nuclear (hyperonic) matter shows that the term $\langle s^{\dagger}s\rangle$ dominates over the other two terms in the odd part of the spectral function, which is responsible for the splitting of the masses between the particle and its antiparticle. The behaviour of the mass splitting at finite densities is similar to the case of strange-charm mesons (as shown in Figs. 5 and 6). Similar to the strange-charm sector, where the isospin asymmetry leads to a drop (rise) of the masses of both $D_{s}^{+}$ and $D_{s}^{-}$ mesons in nuclear (hyperonic) matter, the masses of both $\bar{B_{s}^{0}}$ and ${B_{s}^{0}}$ mesons are also observed to have a downward (upward) shift in nuclear (hyperonic) medium (see figures 11 and 12). However, the mass modifications due to isospin asymmetry are observed to be marginal for T=50 and 100 MeV in hyperonic matter.
For the sake of comparison, we show the density dependence of the mass shifts of the mesons studied in the present work, in the hot nuclear and hyperonic matter in figures 13 and 14 for the charm sector and in figures 15 and 16 for the bottom sector. The masses are plotted for the isospin symmetric ($\eta$=0) as well as for asymmetric matter (with $\eta$=0.5) for values of the temperature, T=50, 100 and 150 MeV. The mass shifts of the open heavy flavor mesons with the non-strange quark (antiquark) constituents are observed to be larger than the strange-charm (strange-bottom) mesons. As has already been mentioned, the antiparticles ($D^{-}$ and $\bar{D^{0}}$ for the charm sector and $B^{+}$ and $B^{0}$ for the bottom sector) have larger mass shifts, as compared to the particles ($D^{+}$ and $D^{0}$ for the charm sector and $B^{-}$ and $\bar{B^{0}}$ for the bottom sector), both in nuclear and hyperonic matter. The same trend of the antiparticle having a larger mass shift as compared to the particle is observed in hyperonic matter, for both strange-charm and strange-bottom mesons, as might be observed in figures 14 and 16, whereas the opposite trend is observed when the strangeness fraction in the medium is zero, i.e. in nuclear medium (see figures 13 and 15). This opposite behaviour of the strange-heavy flavor mesons in nuclear and hyperonic media, as has already been seen from figures 5, 6, 11 and 12, arises due to the term proportional to the strange quark number density in the odd part of the spectral function for hyperonic matter, which is zero in nuclear matter. The isospin asymmetry in the medium leads to a splitting in the masses between the mesons of the iso-doublets ($(D^{0},D^{+})$, $(\bar{D^{0}},D^{-})$ for the open charm mesons and $(B^{-},\bar{B^{0}})$, $(B^{+},B^{0})$ for the open bottom mesons). As may be observed from figures 13 and 14, due to isospin asymmetry in the medium (shown for isospin asymmetry parameter, $\eta$=0.5), there is an upward (downward) shift in the mass of the $D^{-}$ ($\bar{D^{0}}$) for the antiparticle doublet, whereas, there is a drop (rise) in the mass of $D^{+}$ ($D^{0}$) for the particle iso-doublet, in both nuclear and hyperonic matter. The isospin asymmetry is observed to lead to a much larger particle–antiparticle mass difference for the charged sector ($D^{+}$ and $D^{-}$), whereas, there is observed to be a reduction in the particle–antiparticle mass difference for the neutral ($D^{0}$ and $\bar{D^{0}}$) mesons. Hence, the effects of the isospin asymmetry might modify the particle ratios, $D^{+}/D^{-}$ and $D^{0}/\bar{D^{0}}$, as well as, the dilepton yield from the $D\bar{D}$ in asymmetric heavy ion collison experiments. It might be noted here that the results discussed as well as shown in the figures in the present work, correspond to the maximum possible isospin asymmetry ($\eta$=0.5), which is, however, not realized in the heavy ion collision experiments. For a smaller value of isospin asymmetry parameter, $\eta$=0.3 and T=100 MeV, one observes the particle–antiparticle mass difference (in MeV) to be modifed to around 39 and 157 for densities $\rho_{0}$ and $3\rho_{0}$ respectively, for the charged open charm mesons, and, 21 and 83 for the $D^{0}$ and $\bar{D^{0}}$ mesons, from the values of 30 and 121 for the charged as well as neutral open charm mesons in isospin symmetric ($\eta$=0) nuclear matter. The effect due to isospin asymmetry is thus observed to be appreciable, especially at high densities.
The effect of the baryon density is observed to be the dominant medium effect for the modifications of the masses of the open charm and open bottom mesons considered in the present work. There are significant contributions due to the isospin asymmetry of the medium, leading to appreciable increase (decrease) in the particle–antiparticle mass difference in the charged (neutral) mesons at high densities, for both charm ($D$, $\bar{D}$) and bottom ($B$, $\bar{B}$) sectors. With accessible energy for study of charm mesons at the Compressed Baryonic Matter (CBM) Experiment at the Facility for Antiproton and Ion Research (FAIR) at the future facility of GSI TheCBMPhysicsBook ; P_Senger_HIC_FAIR_NICA_Energies , which will be designed to investigate matter at high densites and moderate temperatures, the medium modifications of the masses of the $D$ and $\bar{D}$ mesons could have observable consequences in the particle ratios, e.g., $D^{+}/D^{-}$, $D^{0}/{\bar{D^{0}}}$ in asymmetric nuclear collisions at CBM, FAIR. The medium modifications of the $D$ and $\bar{D}$ mesons can modify the spectral properties of the charmonium states, which can subsequently affect the production of the open and hidden charmed mesons. Also, the dilepton yield arising semi-leptonically from the $D\bar{D}$ correlated pairs can be modified in the isospin asymmetric hadronic matter. The measurement of the invariant dilepton spectra by an electron pair spectrometer at the J-PARC 50-GeV proton spectrometer is being carried out to explore the chiral symmetry of QCD at around nuclear density through study of the medium modifications of the vector mesons produced in pA collsions and decaying to $e^{+}e^{-}$ pairs jparc_16_phi , with the E16 experiment at J-PARC specifically designed to study the mass of the $\phi$ mesons, due to its narrow width. The $K^{*}$ and $K_{1}$ mesons can be produced at J-PARC facility using the kaon beam, and, partial chiral symmetry restoration in the hadronic medium can lead to reduction in the mass difference of these chiral partners. However, due to hadronic rescattering, there is expected to be enhancement in the production ratio of $K^{+}/K^{*}$, which can be a probe for chiral symmetry resoration kstrk1_csb_jparc_16 . In the future, with upgradation of the spectrometer at the J-PARC facility, the measurement of charm hadrons could be possible. The measurements of various hadrons can give a better understanding of the medium modifications of the hadron masses as well as a better insight to chiral symmetry restoration at high densities csb_proposal_Jparc .
@@ V Summary
In the present work, we have studied the medium modifications of the masses of the open charm ($D$, $\bar{D}$, $D_{s}$ and $\bar{D}_{s}$) and open bottom ($B$, $\bar{B}$, $B_{s}$ and $\bar{B}_{s}$) in isospin asymmetric strange hadronic matter at finite temperature. These are computed using a QCD sum rule approach retaining the QCD operators upto dimension 5 in the operator product exapnsion. The values of the QCD operators are calculated within a chiral SU(3) model. The mass splitting of the particles ($D^{0}$, $D^{+}$, $D_{s}^{+}$, $B^{-}$, $\bar{B^{0}}$, $\bar{B_{s}^{0}}$) and their antiparticles ($\bar{D^{0}}$, $D^{-}$, $D_{s}^{-}$, $B^{+}$, ${B^{0}}$, ${B_{s}^{0}}$) arises due to the odd part of the spectral function. In nuclear matter, the masses of the particles, $D^{+}(D^{0})$ and $B^{-}(\bar{B^{0}})$ mesons are observed to be smaller than the masses of the antiparticles, $D^{-}({\bar{D}}^{0})$ and $B^{+}({B^{0}})$. For the strange-charm as well as strange-bottom mesons, the opposite behaviour of the masses of particles ($D_{s}^{+}$ and $\bar{B_{s}^{0}}$) to be higher than the masses of antiparticles ($D_{s}^{-}$ and $B_{s}^{0}$) is observed in nuclear matter, whereas, in the strange hadronic matter, the non-zero value of $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$ leads to the masses of the antiparticles ($D_{s}^{-}$ and $B_{s}^{0}$) to be higher than that of $D_{s}^{+}$ and $\bar{B_{s}^{0}}$ in hyperonic medium. The sign of the difference in the masses of the particle and antiparticle ($m_{+}-m_{-}$) is observed to be opposite in nuclear and hyperonic matter, both for the strange-charm as well as strange-bottom mesons, In the mass splitting between the particle and antiparticle arising due to the odd part of the spectral function, the term proportional to $\langle q_{i}^{\dagger}q_{i}\rangle\equiv\langle s^{\dagger}s\rangle$ for the strange-charm (strange-bottom) mesons thus dominates over the other QCD operators $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$. The effects from density and isospin asymmetry, particularly, on the masses of the $D$, $\bar{D}$ , $B$ and $\bar{B}$ mesons are observed to be appreciable. These can modify e.g., the production of the charmonia and open charm mesons, and, the production ratios $D^{+}/D^{-}$, $D^{0}/\bar{D^{0}}$ in asymmetric relativistic heavy ion collision experiments at Compressed Baryon Matter (CBM) Experiment at the future facility at GSI.